Assuming that the Earth’s population is about 8 billion people: a. How many IPv4 addresses exist per person?
Answer #1
number of bits in an IPv4 address = 32
so, total number of IPv4 addresses = 2^32 = 4294967296

population = 8 billion = 8000000000

number of IPv4 addresses per person = 4294967296/8000000000
= 0.536870912

so, there are only 0.536870912 IPv4 addresses per person on earth.